OpenStudy (anonymous):

what is the factorization of 2x2+18x+36

OpenStudy (uri):

GCF is 2

OpenStudy (uri):

2(x^2+9x+18)

OpenStudy (uri):

Now get a number when we add or subract is 9 and mutiply is 18

OpenStudy (anonymous):

i think its either (x+6)(x+12) or 2(x+6)(x+6)

OpenStudy (uri):

Ah first get the numbers

OpenStudy (anonymous):

im working in the problem but i dont know which of the answers is right

OpenStudy (uri):

Can this be factored more? @joemath314159

OpenStudy (anonymous):

Maybe pulling out the 2 first will make things easier:\[2(x^2+9x+18)\]

OpenStudy (anonymous):

What are two numbers that multiply to 18 and add up to 9?

OpenStudy (anonymous):

3,6

OpenStudy (anonymous):

Right, and since everything is positive, it works out nicely:\[2(x+3)(x+6)\]

OpenStudy (anonymous):

ohhh so it 2(x+6)(x+3)

OpenStudy (mathstudent55):

1. Factor a common factor: \(2x^2+18x+36 \) \(= 2(x^2 + 9x + 18) \) Find two factors of 18 that add to 9: 6 and 3 \(= 2(x + 3)(x + 6) \)

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

thanks

OpenStudy (uri):

lol i said the same thing...up there he/she responded with *ansurs*